# Learning 3D Object Spatial Relationships from Pre-trained 2D Diffusion Models

### v1

@@ Abstract
We present a method for learning 3D spatial relationships between object pairs, referred to as object-object spatial relationships (OOR), by leveraging synthetically generated 3D samples from pre-trained 2D diffusion models. We hypothesize that images synthesized by 2D diffusion models inherently capture plausible and realistic OOR cues, enabling efficient ways to collect a 3D dataset to learn OOR for various unbounded object categories. Our approach begins by synthesizing diverse images that capture plausible OOR cues, which we then uplift into 3D samples. Leveraging our diverse collection of plausible 3D samples for the object pairs, we train a score-based OOR diffusion model to learn the distribution of their relative spatial relationships. Additionally, we extend our pairwise OOR to multi-object OOR by enforcing consistency across pairwise relations and preventing object collisions. Extensive experiments demonstrate the robustness of our method across various object-object spatial relationships, along with its applicability to real-world 3D scene arrangement tasks using the OOR diffusion model.
@@ 1 Introduction
In real-world scenes, specific spatial and functional patterns exist between objects. While some of these patterns are constrained by physical laws (e.g., objects can rest on others but cannot float in mid-air), many originate from functional usage, reflecting how humans interact with and arrange these objects. For instance, chairs are commonly positioned around tables, and items such as cups and bottles are usually placed on tables rather than chairs, despite the physical possibility of alternative arrangements. More sophisticated patterns also exist; for example, a pizza cutter is typically angled relative to the pizza to facilitate slicing across its diameter (as shown in Fig. 1). These intuitive yet diverse relationships between objects, which we refer to as object-object relationships (OOR), capture the relative poses and scales between object pairs. Enabling machines to understand and generate these natural layouts and spatial relationships is crucial for various applications, including content creation, immersive VR/AR experiences, and robotic manipulation tasks aimed at mimicking human behaviors. However, the diversity of object relationships, considering object categories, contexts, and scenarios, makes it challenging to model these patterns through manual annotation or data collection in controlled setups.
@@
One promising approach is to learn spatial relationships between objects from 2D images. However, raw images from the Internet are often cluttered and overly “wild”, complicating the learning of precise 3D spatial relationships between objects. To overcome this limitation, we present an approach to learn 3D object-object relationships from synthetically generated 3D samples capturing plausible OORs. Our method is inspired by recent approaches that pursue human-object interaction and affordances through synthetic images generated by pre-trained image diffusion models [26, 17], where the 2D diffusion model offers effective ways to generate samples for learning such relations.
@@
We begin by rigorously formalizing our definition of OOR, considering the relative scale and pose between object categories. To address the scarcity of available 3D data for learning OORs across diverse object pairs, we introduce a framework that efficiently generates diverse and realistic 3D spatial relationship samples from synthesized 2D images, eliminating the need for labor-intensive data collection or manual annotation. Our 3D lifting pipeline provides a robust solution to reconstruct 3D samples from synthetically generated multi-view 2D images of object-pairs. Next, we present a text-conditioned OOR diffusion based on a score-based model [59, 73] in the spatial parametric space, capable of modeling relative rotations, translations, and scales between objects across diverse scenarios. Our OOR diffusion generates various OOR patterns conditioned on a text prompt describing the scene context. To enhance generalization, we leverage LLM [42] for text context augmentation, expanding our dataset to cover 475 distinct OOR scenarios. In addition, we extend our pairwise OOR modeling to multi-object settings, capturing the spatial relationships among multiple objects within a shared 3D space. Through extensive experiments, we demonstrate the robustness of our method across various object-object spatial relationships. Furthermore, we present several real-world 3D scene arrangement tasks as key applications of our OOR diffusion model.
@@
In summary, our main contributions are as follows: (1) We formulate a novel representation for object-object spatial relationships (OOR); (2) We introduce an effective pipeline for generating diverse 3D OOR data from synthetic images, incorporating prompting strategies, data augmentation, and filtering; (3) We propose a text-conditioned score-based diffusion model to effectively model the OOR distribution. To improve generalization across diverse OOR scenarios, we incorporate LLM-based text augmentation. Moreover, we extend pairwise OOR modeling to the multi-object setting by introducing an optimization strategy guided by novel inference losses; (4) We demonstrate the capability of our OOR diffusion model for 3D scene editing through an optimization technique that directly leverages the score function of the underlying distribution—a key property of score-based models.
@@ 2 Related Work
Learning Object Spatial Relationships. In the field of robotics, research on spatial relations between objects has primarily focused on teaching robots to learn and replicate object placement [52, 24, 45, 37, 40, 20, 29, 67, 3]. However, these methods often overlook complex relationships, like ”a teapot pouring tea into a teacup.” In object detection, there is a growing emphasis on leveraging spatial relationships between objects to improve detection performance, particularly for identifying occluded or challenging-to-recognize objects [76, 33]. For indoor scenes, the availability of 3D datasets [5, 56, 11, 51, 60, 22] capturing object-object spatial relationships has accelerated progress in layout estimation [78, 32], CAD model retrieval [12, 16, 38, 31, 23, 28], and scene generation [44, 70]. However, these methods are limited to predefined object categories, whereas our approach addresses unbounded object pairs.
Learning from Pre-trained 2D Diffusion Models. Advent of diffusion models [55, 18, 7] significantly boosting performance and accelerating advancements in the field. Accordingly, various research efforts have focused on leveraging the vast 2D knowledge embedded in pre-trained diffusion models to extract new insights and enable novel applications. CHORUS [17] and ComA [26] propose a method to learn human-object relationships and interactions from inconsistent 2D image pools generated by diffusion. Our method is inspired by these and models object-object relationships beyond human-object relationships. However, unlike humans, pose estimation of objects is more difficult, making our task a bit more challenging. Moreover, various approaches have been developed to extract semantic features from images using pre-trained diffusion models[75, 69, 71, 72, 39].
Score-based Diffusion Models. A score-based generative model [57, 58] models a distribution using the score function, which is often more tractable than directly computing the likelihood. Also in  [59], it was integrated with diffusion models. Score-based approaches are widely adopted not only for image generation but also for pose generation and estimation tasks for objects [73, 21] and humans [53, 36, 4]. In particular, GenPose [73] is the backbone of our OOR diffusion, which estimates the 6D pose of a given point cloud. However, our OOR diffusion takes text as a condition instead of point clouds, and models not only object poses but also scales.
@@ 3.1 Formulating Object-Object Relationship
We define OOR as the relative poses and scales between a pair of object categories within a context specified by a text prompt $\mathbf{c}$ which describes object categories and their spatial relationships. Each 3D object instance is represented within an instance canonical space where coordinates are defined by centering the origin at the midpoint of the object’s tightest 3D bounding box. The 3D Bbox’s up-vector is aligned to the $y$-axis, and its ground plane is parallel with the $x$-$z$ plane. The frontal side of the object, typically the most observable view, faces the positive $z$-axis, although our method accommodates any canonical orientation. In this canonical space, the scale is normalized such that the longest edge of the 3D bounding box is set to $1$, while keeping the aspect ratios of the objects. See Fig. 2 for an example. To account for varying aspect ratios within the same object category for spatial relation modeling, we further consider the scale-normalized canonical space where the 3D bounding box is normalized into a unit cube with length $1$ with all axes. We consider the scaling factor $\mathbf{s}\in\mathbb{R}^{3}_{+}$ defined per each object instance to recover the original aspect ratio from the scale-normalized one: $\mathbf{x}=\mathbf{s}\cdot\mathbf{\hat{x}}$, where $\mathbf{\hat{x}}\in\mathbb{R}^{3}$ is the 3D point in the scale-normalized canonical space and $\mathbf{x}\in\mathbb{R}^{3}$ is the 3D point in the original instance canonical space with original aspect ratio.
@@
$\mathbf{x}_{\mathcal{T}\rightarrow\mathcal{B}}=\mathbf{R}^{\mathcal{T}\rightarrow\mathcal{B}}\cdot(\mathbf{s}^{\mathcal{T}\rightarrow\mathcal{B}}\odot\mathbf{\hat{x}}_{\mathcal{T}})+\mathbf{t}^{\mathcal{T}\rightarrow\mathcal{B}},$ (2)
where $\odot$ denotes the element-wise product. The $\mathbf{s}^{\mathcal{B}}\in\mathbb{R}^{3}_{+}$ is the scaling factor to convert from the point in the scale-normalized space, $\mathbf{\hat{x}}_{\mathcal{B}}$, into the instance space with the original aspect ratio of it:
@@
Intuitively, an OOR sample $\mathbf{\phi}$ represents a plausible relative spatial relationship for the target object relative to the base object. The use of a non-isotropic scaling factor $\mathbf{s}^{\mathcal{B}}\in\mathbb{R}^{3}_{+}$ that accounts for the original aspect ratio of the 3D BBox is important, as OOR is inherently influenced by object size and shape. For instance, understanding the dimensions and height of a table’s top panel is essential when positioning objects on it, an aspect we approximate using the bounding box scale in our formulation.
@@
To effectively learn OOR defined in Eqs. 1– 3, diverse 3D samples showing plausible spatial relationships of each object pair are required. In particular, we consider the distributions conditioned on each text prompt $\mathbf{c}$, to capture the contextual nuances of these relationships, as spatial configurations can vary across different contexts even for the same object pair. Given the fact that obtaining such 3D samples from the real-world is challenging, we present a novel pipeline for synthetically collecting diverse 3D samples, leveraging pre-trained 2D diffusion models and an advanced 3D uplifting process to generate 3D OOR samples.
@@
High-quality 2D OOR Images Synthesis. We use an off-the-shelf text-to-image model [2] to generate images that aligned to the desired OOR context specified by a text prompt $\mathbf{c}$. We design the prompt with specific strategies to facilitate the later 3D lifting process, including: (1) appending “white background” to the end of the prompt to ensure the full shape of each object is clearly visible; (2) incorporating object shape descriptions to better align with our template object meshes for categories with diverse shapes, such as chairs; (3) Adjusting the viewpoint to mitigate image frame constraints for category pairs with large scale differences, such as (table, teacup). Additionally, we further diversify the OOR image samples using the image-to-video model [62] to generate diverse OORs within the context. Each frame of the synthesized video is used as an additional synthesized 2D sample, ignoring temporal information. Further discussion of prompting and augmentation strategies is provided in Sec. A.
@@
Pseudo Multi-view Generation and SfM. Given an image that includes the OOR cues for the target object pair, we produce pseudo-multi-view images using an off-the-shelf novel view synthesis method, SV3D [64], which generates novel view images on the orbital viewpoint changes by taking a single image input. To enhance multi-view synthesis quality, we first apply a segmentation method [50, 30, 25, 35] to remain the relevant object pair only, removing all background clutter. While the pseudo-multi-view synthesis produces visually plausible outputs, it does not guarantee perfect geometric consistency across views. Interestingly, we find that the recent learning-based structure-from-motion (SfM) approach, VGGSfM [65], produces plausible 3D reconstruction outputs from synthesized multi-view inputs. However, this method still occasionally fails to produce reliable 3D outputs, so we filter out unsuccessful reconstructions by checking the number of 3D points in the resulting point clouds. As the output of SfM, we obtain the 3D point cloud $\mathcal{P}=\{\mathbf{P}_{j}\}_{j=1}^{N},\;\mathbf{P}_{j}\in\mathbb{R}^{3}$, and their corresponding 2D keypoints, $\{\mathbf{p}_{j}^{k}\}_{k=1}^{m_{j}},\;\mathbf{p}_{j}^{k}\in\mathbb{R}^{2}$, where $N$ denotes the number of 3D points, and $m_{j}$ is the number of corresponding 2D keypoints for the $j$-th 3D point.
@@
Pose and Scale Extraction through Mesh Registration. Given the point cloud $\mathcal{P}$ reconstructed in a certain coordinate system (denoted as point cloud space), we register the base object’s template mesh $\mathcal{M}_{\mathcal{B}}$ and the target object’s template mesh $\mathcal{M}_{\mathcal{T}}$ from their canonical spaces to the point cloud. From this registration, we compute their relative spatial transformation $T_{\text{rel}}$, which is parameterized by Eq. 1.
We first separate the 3D points cloud into the base object part and the target object part, resulting in the base object points $\mathcal{P}_{\mathcal{B}}$ and target object points $\mathcal{P}_{\mathcal{T}}$. To achieve this, we apply a 2D segmentation method on the pseudo-multi-view 2D images. We find that treating the pseudo-multi-view images as a video and applying a video segmentation model [48, 49] shows better segmentation quality. Next, we register each template mesh into the corresponding point cloud. For simplicity, we omit the object indicator, and denote the point clouds as $\mathcal{P}$ and the template mesh in the canonical space as $\mathcal{M}$.
@@
To register each mesh to the point cloud space, we leverage the semantic correspondence features [72] by extracting the cues from 2D views. Specifically, for each 3D point $\mathbf{P}_{j}$, we consider its corresponding 2d projections in the corresponding 2D image space $\{\mathbf{p}_{j}^{k}\}_{k=1}^{m_{j}}$ and extract the semantic features $\mathbf{f}_{j}^{k}\in\mathbb{R}^{F}$ at each pixel via the method of [72], where $F$ is the semantic feature dimension ($F$=768 in practice). We apply the same 2D semantic feature extraction process for each point on the template mesh model $\mathbf{M}_{j}$ from its rendered 2D views. Then we aggregate the semantic features from 2D views into a single feature vector for the corresponding 3D point. Rather than directly taking the mean of the 2D feature vectors, we find that it is more effective to reduce the dimension of the semantic features by applying PCA [19] as a way to extract more important signals and reduce noisy parts. To apply PCA, we aggregate all 2D features from both pseudo-multi-view images and rendered template mesh images, and reduce the semantic feature dimensions to ${\mathbf{f}^{\prime}}_{j}^{k}\in\mathbb{R}^{F^{\prime}}$, where ${F^{\prime}}<F$ ($F^{\prime}$= 15 in practice). Finally, for each 3D point, we compute the mean feature vector from the corresponding 2D features in the reduced feature dimension. The effectiveness of this PCA-based dimensionality reduction for registration is demonstrated in our experiments (Sec.  4.4).
@@
Given the semantic features associated with the 3D point cloud and template mesh points, we establish correspondences by identifying nearest neighbors based on cosine similarity among semantic features. We retain only those whose similarity exceeds a certain threshold. We use Procrustes analysis [15, 14] with RANSAC [10] for the registration, followed by ICP to further refine it [1]. To this end, we obtain the transformation $T$ from $\mathcal{M}$ to $\mathcal{P}$:
@@
where $T$ is parameterized by 3D rotation $\mathbf{R}\in SO(3)$, 3D translation $\mathbf{t}\in\mathbb{R}^{3}$, and an isotropic scaling factor $s\in\mathbb{R}_{+}$, transforming the point in the mesh template $\mathcal{M}$ into the point cloud space $\mathcal{P}$
By applying this registration process per each template model, we can compute the transformation from each mesh template model into a common point cloud space: $T_{\mathcal{B}\rightarrow\mathcal{P}}$ and $T_{\mathcal{T}\rightarrow\mathcal{P}}$. Finally, we can find the relative pose and scale of the $\mathcal{M}_{\mathcal{T}}$ relative to $\mathcal{M}_{\mathcal{B}}$, denoted as $T_{\mathcal{T}\rightarrow\mathcal{B}}$:
@@
where $T_{\mathcal{P}\rightarrow\mathcal{B}}$ is the inverse transformation of $T_{\mathcal{B}\rightarrow\mathcal{P}}$. Note that we use an isotropic scaling factor $s\in\mathbb{R}_{+}$ for simplicity during the registration by keeping the original aspect ratio. After the registration, we multiply the size of the tightest 3D bounding box to represent the same 3D bbox scale representation defined in Eq. 1, as follows:
$\begin{gathered}\mathbf{s}^{\mathcal{T}\rightarrow\mathcal{B}}:=s_{\mathcal{T}\rightarrow\mathcal{B}}*\text{BBOX}(\mathcal{T}),\\ \mathbf{s}^{\mathcal{B}}:=\text{BBOX}(\mathcal{B}).\end{gathered}$ (6)
Fig. 3 shows the overview of OOR dataset generation process. In practice, to account for the shape difference between the template mesh and the reconstructed outputs, we use several template meshes as candidates and select the one that best fits the reconstructed object based on the DINO [43, 6] features. We also apply a series of automatic filtering processes to remove unreliable samples. See Sec. A for details.

### v2

@@ Abstract
We present a method for learning 3D spatial relationships between object pairs, referred to as object-object spatial relationships (OOR), by leveraging synthetically generated 3D samples from pre-trained 2D diffusion models. We hypothesize that images synthesized by 2D diffusion models inherently capture realistic OOR cues, enabling efficient collection of a 3D dataset to learn OOR for various unbounded object categories. Our approach synthesizes diverse images that capture plausible OOR cues, which we then uplift into 3D samples. Leveraging our diverse collection of 3D samples for the object pairs, we train a score-based OOR diffusion model to learn the distribution of their relative spatial relationships. Additionally, we extend our pairwise OOR to multi-object OOR by enforcing consistency across pairwise relations and preventing object collisions. Extensive experiments demonstrate the robustness of our method across various object-object spatial relationships, along with its applicability to 3D scene arrangement tasks and human motion synthesis using our OOR diffusion model.
@@ 1 Introduction
In real-world scenes, specific spatial and functional patterns exist between objects. While some are constrained by physical laws (e.g., objects can rest on others but cannot float in mid-air), many arise from functional usage, reflecting how humans interact with and arrange these objects. For instance, chairs are commonly positioned around tables, and items such as cups and bottles are usually placed on tables rather than chairs, despite the physical possibility of alternative arrangements. More sophisticated patterns also exist; for example, a pizza cutter is typically angled relative to the pizza to facilitate slicing across its diameter (as shown in Fig. 1). These intuitive yet diverse relationships, which we refer to as object-object spatial relationships (OOR), capture the relative poses and scales between object pairs. Enabling machines to understand and generate these natural layouts and spatial relationships is crucial for various applications such as content creation, immersive VR/AR, and robotic manipulation tasks. However, the diversity of object spatial relationships, across object categories, contexts, and scenarios, makes them challenging to model through manual annotation or data collection in controlled setups.
@@
One promising approach is to learn spatial relationships between objects from 2D images. However, raw images from the Internet are often cluttered and overly “wild”, complicating the learning of precise 3D spatial relationships between objects. To overcome this limitation, we present an approach to learn 3D object spatial relationships from synthetically generated 3D samples capturing plausible OORs. Our method is inspired by recent approaches that pursue human-object interaction and affordances through synthetic images generated by pre-trained image diffusion models [26, 18], where the 2D diffusion model offers effective ways to generate samples for learning such relations.
@@
We first rigorously formalize OOR as the relative scale and pose between two object categories. To address the scarcity of 3D data for learning OORs across diverse object pairs, we introduce a framework that generates diverse and realistic 3D spatial relationship samples from synthesized 2D images, eliminating the need for labor-intensive data collection or manual annotation. Our 3D lifting pipeline robustly reconstructs 3D samples from synthetically generated 2D images of object-pairs. Next, we present a text-conditioned OOR diffusion based on a score-based model [58, 72] in the spatial parametric space, capable of modeling relative rotations, translations, and scales between objects across diverse scenarios. Our OOR diffusion generates varied OOR patterns conditioned on a text prompt describing the scene context. To enhance generalization, we leverage LLM [41] for text context augmentation, expanding our dataset to 475 distinct OOR scenarios. In addition, we extend our pairwise OOR modeling to multi-object settings, capturing spatial relationships among multiple objects in a shared 3D space. Through extensive experiments, we demonstrate the robustness of our method across various object-object spatial relationships. We also present several 3D scene arrangement tasks and the synthesis of human motion interacting with two objects as applications of our OOR diffusion model.
@@
In summary, our main contributions are as follows: (1) We formulate a novel representation for object-object spatial relationships (OOR); (2) We introduce an effective pipeline to generate diverse 3D OOR data from synthetic images, incorporating prompting strategies, data augmentation, and filtering; (3) We propose a text-conditioned score-based diffusion model to effectively model the OOR distribution. To improve generalization across diverse OOR scenarios, we incorporate LLM-based text augmentation. We also extend pairwise OOR modeling to the multi-object setting via an optimization strategy guided by novel inference-time losses; (4) We demonstrate the capability of our OOR diffusion model for 3D scene editing through optimization that directly leverages the score function of the underlying distribution. We further show that the generated OOR can be used to synthesize human motions interacting with two objects.
@@ 2 Related Work
Learning Object Spatial Relationships. In robotics, research on spatial relations between objects has primarily focused on teaching robots to learn and replicate object placement [51, 25, 45, 36, 39, 21, 30, 66, 4]. However, these methods often overlook complex relationships, like ”a teapot pouring tea into a teacup.” In object detection, there is a growing emphasis on leveraging spatial relationships between objects to improve detection performance, particularly for identifying occluded or challenging-to-recognize objects [75, 33]. For indoor scenes, the availability of 3D datasets [6, 55, 12, 50, 59, 23] capturing object-object spatial relationships has accelerated progress in layout estimation [77, 32], CAD model retrieval [13, 17, 37, 31, 24, 29], and scene generation [43, 69]. However, these methods are limited to predefined object categories, whereas our approach addresses unbounded object pairs.
Learning from Pre-trained 2D Diffusion Models. The advent of diffusion models [54, 19, 8] has significantly boosted performance and accelerated progress in the field. Accordingly, many studies have focused on leveraging the rich 2D knowledge in pre-trained diffusion models to extract insights and enable novel applications. CHORUS [18] and ComA [26] propose learning human-object relationships and interactions from inconsistent 2D image pools generated by diffusion. However, unlike humans, estimating object poses is more difficult, making our task more challenging. Additionally, various approaches have been developed to extract semantic features from images using pre-trained diffusion models [74, 68, 70, 71, 38, 3].
Score-based Diffusion Models. A score-based generative model [56, 57] models a distribution using the score function, often more tractable than computing the likelihood directly. It was integrated with diffusion models in [58]. Score-based methods are widely adopted not only for image generation but also for pose generation and estimation of objects [72, 22] and humans [52, 35, 5]. In particular, GenPose [72], which estimates the 6D pose of a point cloud, serves as the backbone of our OOR diffusion. DAViD [27] applies score-based modeling to the HOI domain by leveraging GenPose.
@@ 3.1 Formulating Object-Object Relationship
We define OOR as the relative poses and scales between a pair of object categories specified by a text prompt $\mathbf{c}$ describing object categories and their spatial relationships. Each 3D object instance is placed in an instance canonical space whose origin is the midpoint of its tightest 3D bounding box. The 3D Bbox’s up-vector aligns with the $y$-axis, and its ground plane is parallel with the $x$-$z$ plane. The frontal side, typically the most observable view, faces the $z$-axis, although our method accommodates any canonical orientation. The scale is normalized such that the longest edge of the 3D Bbox is set to $1$, while keeping the aspect ratios of the objects. See Fig. 2 for an example. To account for varying aspect ratios within a category, we further consider a scale-normalized canonical space that normalizes the 3D Bbox to a unit cube. We consider the per-instance scaling factor $\mathbf{s}\in\mathbb{R}^{3}_{+}$ which recover the original aspect ratio from the scale-normalized one: $\mathbf{x}=\mathbf{s}\odot\mathbf{\hat{x}}$, where $\mathbf{\hat{x}}\in\mathbb{R}^{3}$ is the 3D point in the scale-normalized canonical space, $\mathbf{x}\in\mathbb{R}^{3}$ is the 3D point in the original instance canonical space with original aspect ratio, and $\odot$ denotes the element-wise product.
@@
$\mathbf{x}_{\mathcal{T}\rightarrow\mathcal{B}}=\mathbf{R}^{\mathcal{T}\rightarrow\mathcal{B}}\cdot(\mathbf{s}^{\mathcal{T}\rightarrow\mathcal{B}}\odot\mathbf{\hat{x}}_{\mathcal{T}})+\mathbf{t}^{\mathcal{T}\rightarrow\mathcal{B}}.$ (2)
The $\mathbf{s}^{\mathcal{B}}\in\mathbb{R}^{3}_{+}$ is the scaling factor to convert from the point in the scale-normalized space, $\mathbf{\hat{x}}_{\mathcal{B}}$, into the instance space with the original aspect ratio of it:
@@
Intuitively, an OOR sample $\mathbf{\phi}$ represents a plausible relative spatial relationship for the target object relative to the base object. The use of a non-isotropic scaling factor $\mathbf{s}^{\mathcal{B}}\in\mathbb{R}^{3}_{+}$ that accounts for the original aspect ratio of the 3D BBox is important, as OOR is inherently influenced by object 3D scale. For instance, placing items on a table requires knowing the dimensions of the tabletop and height, which we approximate with its 3D Bbox scale.
@@
To effectively learn OOR defined in Eqs. 1– 3, diverse 3D samples with plausible spatial relationships of each object pair are required. Obtaining such 3D samples from the real world is challenging. We present a novel pipeline that synthesizes diverse 3D samples by leveraging pre-trained 2D diffusion models and an advanced 3D uplifting process.
@@
High-quality 2D OOR Images Synthesis. We use an off-the-shelf text-to-image model [2] to generate images that are aligned to the OOR context in text prompt $\mathbf{c}$. We design the prompt with specific strategies to facilitate the later 3D lifting process, including: (1) appending “white background” to the end of the prompt to ensure the full shape of each object is clearly visible; (2) incorporating object shape and texture descriptions to better align with our template object meshes; (3) Adjusting the viewpoint to mitigate image frame constraints for category pairs with large scale differences, such as (table, teacup). Additionally, we further diversify OORs by passing each image to an image-to-video model [61] and treating every frame as an additional 2D sample, ignoring temporal order. Further discussion of prompting and augmentation strategies is provided in Sec. A
@@
Pseudo Multi-view Generation and SfM. Given an image containing the OOR cues for the object pair, we produce pseudo-multi-view images using an off-the-shelf novel view synthesis method, SV3D [63], which synthesizes circular multi-views from a single image input. These pseudo-multi-views look visually plausible often but may lack geometric consistency across views, so we reconstruct them with VGGSfM [64] and discard failures. As the output of SfM, we obtain the 3D point cloud $\mathcal{P}=\{\mathbf{P}_{j}\}_{j=1}^{N},\;\mathbf{P}_{j}\in\mathbb{R}^{3}$, and their corresponding 2D keypoints, $\{\mathbf{p}_{j}^{k}\}_{k=1}^{m_{j}},\;\mathbf{p}_{j}^{k}\in\mathbb{R}^{2}$, where $N$ denotes the number of 3D points, and $m_{j}$ is the number of corresponding 2D keypoints for the $j$-th 3D point.
@@
Pose and Scale Extraction through Mesh Registration. Given the point clouds $\mathcal{P}$ reconstructed in a certain coordinate system (denoted as point cloud space), we register the base object’s template mesh $\mathcal{M}_{\mathcal{B}}$ and the target object’s template mesh $\mathcal{M}_{\mathcal{T}}$ from their canonical spaces to the point cloud. From this registration, we compute their relative spatial transformation $T_{\text{rel}}$, which is parameterized by Eq. 1.
@@
We first partition $\mathcal{P}$ into base object points $\mathcal{P}_{\mathcal{B}}$ and target object points $\mathcal{P}_{\mathcal{T}}$. Treating pseudo-multi-view images as a video and applying a video-segmentation model [48, 49] yields better segmentation quality. Next, we register each template mesh into the corresponding point cloud. For simplicity, we omit the object indicator ($\mathcal{B}$ or $\mathcal{T}$), and denote the point clouds as $\mathcal{P}$ and the template mesh in the canonical space as $\mathcal{M}$. To register each mesh, we extract semantic correspondence features [71] from 2D views. Specifically, for each 3D point $\mathbf{P}_{j}$, its corresponding image projections $\{\mathbf{p}_{j}^{k}\}_{k=1}^{m_{j}}$ yield features $\mathbf{f}_{j}^{k}\in\mathbb{R}^{F}$ at each pixel, where $F$ is the semantic feature dimension (FF=768 in practice). We apply the same process for each point on $\mathcal{M}$ from its rendered 2D views. Then we aggregate the semantic features from 2D views into a single feature vector for the corresponding 3D point. Rather than directly averaging these 2D features, we first run PCA [20] on all (pseudo-multi-view and rendered) features to obtain ${\mathbf{f}^{\prime}}_{j}^{k}\in\mathbb{R}^{F^{\prime}}$, where ${F^{\prime}}<F$ ($F^{\prime}$= 15 in practice). Finally, average them per 3D point.
@@
Using the semantic features of $\mathcal{P}$ and $\mathcal{M}$, we establish correspondences by identifying nearest neighbors based on cosine similarity. We retain only those whose similarity exceeds a certain threshold. We use Procrustes analysis [16, 15] with RANSAC [11] for the registration, followed by ICP [1] to further refine it. To this end, we obtain the transformation $T$ from $\mathcal{M}$ to $\mathcal{P}$:
@@
where $T$ is parameterized by 3D rotation $\mathbf{R}\in SO(3)$, 3D translation $\mathbf{t}\in\mathbb{R}^{3}$, and an isotropic scaling factor $s\in\mathbb{R}_{+}$, transforming the point in the $\mathcal{M}$ into the point cloud space.
Applying the registration to each template mesh gives $T_{\mathcal{B}\rightarrow\mathcal{P}}$ and $T_{\mathcal{T}\rightarrow\mathcal{P}}$, the transforms from the base and target meshes to the point cloud space. Finally, we can find the relative pose and scale of the $\mathcal{M}_{\mathcal{T}}$ relative to $\mathcal{M}_{\mathcal{B}}$, denoted as $T_{\mathcal{T}\rightarrow\mathcal{B}}$:
@@
where $T_{\mathcal{P}\rightarrow\mathcal{B}}$ is the inverse transformation of $T_{\mathcal{B}\rightarrow\mathcal{P}}$. Note that we use an isotropic scale $s\in\mathbb{R}_{+}$ during the registration by keeping the original aspect ratio. After the registration, we obtain the scale representation defined in Eq. 1, as follows:
$\begin{gathered}\mathbf{s}^{\mathcal{T}\rightarrow\mathcal{B}}:=s_{\mathcal{T}\rightarrow\mathcal{B}}\cdot\text{BBOX}(\mathcal{T}),\\ \mathbf{s}^{\mathcal{B}}:=\text{BBOX}(\mathcal{B}).\end{gathered}$ (6)
Fig. 3 overviews the OOR dataset generation process. To account for the shape deviations, we use several template meshes as candidates and select the best via DINO features [42, 7]. Further, we apply a series of automatic filtering processes to remove unreliable samples. See Sec. A for details.